# Model Inversion Attacks Meet Cryptographic Fuzzy Extractors

### v1

@@ 1 Introduction
Machine learning (ML) systems are increasingly being used in security-sensitive applications, such as for authenticating users [1, 2], establishing content ownership [3], CSAM detection [3], and so on. In these applications, neural networks often encode security-sensitive inputs into vector representations called embedding vectors. For example, in face authentication, standard neural networks generate embeddings corresponding to a user’s face image [1, 2]. These embedding vectors are stored on databases of application service providers [4] or on authentication devices [5].
Databases and devices storing sensitive information are frequently prone to breaches [6] and theft/loss [5] respectively. This gives rise to a threat model where the attacker gains access to all information persistently stored in the database or the device, as is typical of attacks colloquially referred to “smash-and-grab” attacks. The breached information can be used to recover the original security-sensitive data corresponding to the embedding vectors, causing severe privacy risks. A well-known class of such attacks is called model inversion. Since their discovery in 2015 on deep networks [7], no defenses with strong formal security have yet been developed for model inversion attacks.
Model inversion attacks allow the adversary to recover inputs, given the outputs of the ML model. The canonical example introduced by the original work of Fredrikson et al. showcases the threat to face recognition systems, where outputs of deep networks can be used to recover the input face images reasonably accurately [7]. Over the last decade, ML models have evolved to serve tasks beyond classification which motivated the original work of Fredrikson et al. The outputs of modern ML models are embedding vectors which can be compared in $\ell_{2}$ distance to other vectors. These vectors can be used to implement classification in a closed-set [8], similarity testing in an open-set [9, 2], lookup in RAG databases [10], or generative tasks like text-to-image generation [11]. Recent works have shown that embedding vectors generated from today’s ML models, if stored unprotected (in the clear), are highly susceptible to inversion attacks [12, 13, 14]. Therefore, it is important to devise a generic defense for model inversion attacks that works for applications where vector comparison with the standard Euclidean ($\ell_{2}$) distance is used. In this paper, we will use the classical example of face recognition ML models, but our goal is to motivate a broader defense strategy.
@@
One way to build a generic defense with provable guarantees is to turn to cryptographic methods. Specifically, can one post-process the output of the ML model with a cryptographic hash [^1] One could encrypt the embeddings with standard encryption. But, then the encryption key itself is persistently stored and offers no security when all persistent data is leaked. before they are stored, to create what we call protected embeddings, such that they can still be used for $\ell_{2}$-based authentication while resisting model inversion attacks? The advantage of post-processing defenses is that they do not require re-training of the ML model, and they are agnostic to any specific model inversion attack. They offer a complementary approach to many in-training defense methods [15, 16, 17, 18], which offer no formal security guarantees, nor do they aim to develop an attack-agnostic defense. Remarkably, despite their appeal, post-processing defenses have not yet been considered for tackling model inversion.
To understand why, it is perhaps useful to draw parallels with how hashing protects against the same threat model of password database breaches [19]. It is standard practice to store passwords salted-and-hashed using standard cryptographic hash functions like SHA256, so that inversion of the plaintext passwords from a leaked password database becomes hard [20]. However, such hash functions do not work for protecting embeddings generated from ML models. This is because ML models are designed to tolerate noise: Embeddings generated from two different face images are considered to be of the same person if their generated embedding vectors are close in $\ell_{2}$ distance. In face authentication, the user being authenticated must have an embedding vector that is close to the one expected, unlike passwords that must match exactly. However, standard cryptographic hash functions (eg. SHA256) used for salting-and-hashing passwords cannot tolerate any errors in inputs, i.e., even a single bit change in input induces a large change in output of the hash function. Hence, standard hash functions like SHA256 cannot be used to protect ML embedding vectors.
@@
Post-processing defenses for tackling model inversion attacks on modern ML models have not been demonstrated so far. Nonetheless, there have been partially-secure post-processing candidates that one could leverage from prior literature as a starting point [21, 22]. These schemes allow noisy inputs to authenticate successfully as desired. Indeed, in our evaluation, we show that these schemes are effective against prior model inversion attacks which target unprotected embeddings [14], reducing their success rate from 99% to $2\%$. However, as we will show, prior protection schemes are still breakable and do not enjoy cryptographically strong properties that are ideally desired.
@@
We evaluate our Pipe attack against two of the most potent partial protection mechanisms we know of: Facial-Fuzzy Extractor (Facial-FE) [23] and Multispace Random Projection (MRP) [24]. Pipe highlights the weaknesses of these constructions when applied to face authentication under the realistic threat model of full data leakage. Experimentally, Pipe has attack success rate of over 47% for embeddings generated from state-of-the-art models like Facenet [2] and ArcFace [1] against both of these partial defenses (see Tables III and IV). In most settings, the observed attack success rate is well over 89%.
Figure 1 shows a successfully recovered face from Facial-FE-protected embeddings using Pipe for a quick illustration. It is close to the original image both visually and as per image similarity metrics like LPIPS [9], showing the privacy-invasive nature of the attack. The normalised embeddings of these images are also close in $\ell_{2}$ distance.
@@
Does there exist a cryptographic primitive for $\ell_{2}$-based authentication that would provably defeat all model inversion attacks, regardless of the underlying ML model, with the full-leakage threat model? Three properties would be required by such a primitive: noise tolerance, security under full leakage, and utility. One can think of it as an analogue to the salt-and-hash mechanism used for passwords, except that this primitive must tolerate noise in the input vector. Additionally, it should be hard to recover an input close to the original input even when given all outputs. We introduce this ideal primitive in Section 3 and present its formal definition in appendix A.
As a key conceptual contribution (see Theorem 2), we show that the ideal primitive can be achieved if one constructs a secure fuzzy extractor—a cryptographic construct [25] that predates model inversion attacks by a decade—for the desired $\ell_{2}$ norm vector comparison. Finally, as a technical contribution, we propose a candidate construction called the L2FE-Hash, a lattice-based fuzzy extractor (FE) that enables $\ell_{2}$ norm vector comparison on hashed outputs. We formally show that it is computationally secure under certain assumptions on the input distribution (see Theorem 1). We empirically show that it has reasonable accuracy for face authentication with state-of-the-art ML models. We also confirm that L2FE-Hash resists prominent prior model inversion attacks as well as our adaptive Pipe attack in our experimental evaluation (see Section 8).
@@
While we have chosen the face authentication system as a concrete application, aligning with the original work on model inversion [7], the threat of inversion attacks on embedding vectors extends well beyond that. All our security properties, attacks, and definitions are generic—they are defined abstractly for models that produce similar high-dimensional embedding vectors for similar inputs. We define closeness with respect to $\ell_{2}$ distance since it is common in face authentication, but one can easily extend our definitions to any distance metric. One can readily consider modalities that rely on $\ell_{2}$ distance, like audio and text [26, 27] in future work. Likewise, one can consider other security applications such as image similarity checking for copyright infringement or CSAM detection [3].
@@
To the best of our knowledge, this is the first work that connects two long-standing but disparate lines of research, namely on model inversion attacks and on cryptographic fuzzy extractors. First, we show a new model inversion attack called Pipe which reliably defeats existing post-processing mechanisms, highlighting the lack of practical defenses with desirable cryptographic properties. Second, we describe a candidate fulfilling all ideal properties practically, called L2FE-Hash. We prove that fuzzy extractors are robust against inversion attacks generically, and show that L2FE-Hash is an $\ell_{2}$-norm fuzzy extractor construction that is empirically robust against Pipe, while having a usable accuracy for face authentication systems.
@@ 2 The Problem
Protecting long-term secrets, like authentication credentials, is an important goal in the event of catastrophic data breaches. We explain how this problem manifests in the context of vector embeddings used in modern ML models for face authentication, following the inspiration set in [7].
@@ 2.1 Application: ML Face Authentication
Billions of users have had password credentials leaked via data breaches, posing serious threats to security and privacy. As authentication services, for example ML-based face authentication like ClearView [4] and crypto wallets [5] store credentials (embeddings) derived from face images persistently, they face a similar risk from breach or theft [6].
@@ 2.2 Inversion of Unprotected ML Embeddings
Consider an attacker who gets unauthorized access to the database storing sensitive embedding vectors persistently. If no protection mechanism was implemented (i.e. $\mathcal{M}$ outputs were stored directly in Figure 2), then the “unprotected” embedding $x$ retains significant information about the original image, and there is a risk that the original image can be inverted from $x$ approximately. Recent prior works have shown that ML models can be trained [12, 13, 14] to perform such inversion for standard face recognition ML systems ($\mathcal{M}$) such as Facenet [2] and ArcFace [1]. If successful, the inverted image is close enough to the original face image that it passes authentication with $\mathcal{M}$. Prior works have reported attack success rates of over $90\%$ (see Section 6 too). Thus, highly effective inversion attacks on unprotected embeddings are known, posing a serious privacy risk.
@@ 2.3 Protecting Embeddings via Post-processing
A good defense mechanism should protect embedding vectors derived from sensitive inputs when they are persistently stored. Ideally, the defense should not be attack-specific and should generically defeat all feasible attacks. Moreover, we are not seeking solutions for securing persistent database storage via additional trust in hardware [28] or in third parties to securely store cryptographic keys.
@@
Many prior defenses are not in the post-processing regime. They are methods that work during model training to reduce the amount of sensitive information held by the resulting embeddings [15, 18, 16, 17]. But such mechanisms have had limited success. Recent works show that the defended model remains susceptible to prior inversion attacks, with $\geq 30\%$ users being re-identified for models that have accuracy $\geq 80\%$ [15, 17, 29]. Therefore, we leave this alternative line of mechanisms out of scope. Besides, they are complementary to post-processing mechanisms.
@@ 2.4 Threat Model
To keep attack assumptions minimal, we work in the black-box threat model. We adapt the black-box model presented in [7] and assume the attacker only has query (or API) access to the target model $\mathcal{M}_{prot}$ (see Figure 2). The query exposes protected embeddings, instead of classification results as presumed in [7] for generality, since ML models today are used in non-classification tasks too. Moreover, protection schemes used by the target models are publicly available algorithms known to the attacker.
We assume that the attacker has access to a public dataset of face images, similar to that assumed in the original attack [7]. Our public dataset does not overlap with the leaked database, and the attacker can query $\mathcal{M}_{prot}$ on this dataset to obtain corresponding embeddings. These assumptions are practical and common to prior attacks [14, 30, 31]. This is also analogous to how password crackers use real-world password distributions gleaned from leaked password databases [19, 32].
@@ 3 An Ideal Post-Processing Primitive
The need for post-processing protection mechanisms has been recognized in prior cryptographic literature [21, 22], but their desired security properties have not been carefully established in the context of model inversion. In particular, no secure mechanisms from prior works satisfy the desired properties, as we will experimentally show in Section 6.
@@ II Input image obtained during enrollment
Item 3 is necessary to prevent trivial functions that compress inputs into a very small output domain. If output entropy is too small, the adversary would be able to guess inputs by brute-forcing. Moreover, it would render the primitive useless in applications that want to use the output to distinguish various inputs. For example, a primitive designed for face authentication should map input embeddings corresponding to different users to different protected embeddings. Hence, the output entropy required for Item 3 depends on the application use case and is often related to the input entropy. We use the standard computational HILL entropy here, please see Definition 3 in Appendix A.
@@ 3.2 Relation to Fuzzy Extractors
We will later show formally that the desired properties are achieved by a fuzzy extractor. Briefly, the Gen and Rep subroutines of a fuzzy extractor work similarly to the Gen and Rep subroutines of our ideal primitive, albeit with different security formulations. However, existing fuzzy extractor constructions are designed for $\ell_{0}$ distances, such as the Hamming distance or set difference, to suit the use case of iris or fingerprint authentication [25]. There is no secure construction of an $\ell_{2}$-based fuzzy extractor, as needed for practical use with face distributions yet. We address this gap by proposing an ideal primitive for an $\ell_{2}$-based authentication system, and give a working candidate construction L2FE-Hash of the ideal primitive in Section 7.
@@ 4 Existing Protection Schemes
As a starting point for post-processing defenses, one can consider two cryptographic schemes that are state-of-the-art of their respective family of concepts: a Facial-Fuzzy Extractor and Multispace Random Projection. We detail their constructions and explain why they only partially-secure under the threat model of full leakage. They serve as two main targets for our new attack Pipe presented in Section 5, and experimentally confirm it in Section 6.

### v2

@@ 1 Introduction
Machine learning (ML) systems are increasingly being used in security-sensitive applications, such as for authenticating users [2022:arcface, 2015:facenet], establishing content ownership [2021:neural-hash], CSAM detection [2021:neural-hash], and so on. In these applications, neural networks often encode security-sensitive inputs into vector representations called embedding vectors. For example, in face authentication, standard neural networks generate embeddings corresponding to a user’s face image [2022:arcface, 2015:facenet]. These embedding vectors are stored on databases of application service providers [2023:clearview-ai] or on authentication devices [2024:biometric-wallet].
Databases and devices storing sensitive information are frequently prone to breaches [2024:outabox-breach] and theft/loss [2024:biometric-wallet] respectively. This gives rise to a threat model where the attacker gains access to all information persistently stored in the database or the device, as is typical of attacks colloquially referred to “smash-and-grab” attacks. The breached information can be used to recover the original security-sensitive data corresponding to the embedding vectors, causing severe privacy risks. A well-known class of such attacks is called model inversion. Since their discovery in 2015 on deep networks [2015:model-inversion], no defenses with strong formal security have yet been developed for model inversion attacks.
Model inversion attacks allow the adversary to recover inputs, given the outputs of the ML model. The canonical example introduced by the original work of Fredrikson et al. showcases the threat to face recognition systems, where outputs of deep networks can be used to recover the input face images reasonably accurately [2015:model-inversion]. Over the last decade, ML models have evolved to serve tasks beyond classification which motivated the original work of Fredrikson et al. The outputs of modern ML models are embedding vectors which can be compared in $\ell_{2}$ distance to other vectors. These vectors can be used to implement classification in a closed-set [2021:text-embedding-classification], similarity testing in an open-set [2018:LPIPS, 2015:facenet], lookup in RAG databases [2024:rag-databases], or generative tasks like text-to-image generation [2021:text-to-image]. Recent works have shown that embedding vectors generated from today’s ML models, if stored unprotected (in the clear), are highly susceptible to inversion attacks [2023:stylegan-recontruction-attack, 2023:id3pm-diffusion-inversion-attack, 2024:bob]. Therefore, it is important to devise a generic defense for model inversion attacks that works for applications where vector comparison with the standard Euclidean ($\ell_{2}$) distance is used. In this paper, we will use the classical example of face recognition ML models, but our goal is to motivate a broader defense strategy.
@@
One way to build a generic defense with provable guarantees is to turn to cryptographic methods. Specifically, can one post-process the output of the ML model with a cryptographic hash [^1] If we encrypt the embeddings instead, the encryption key must be persistently stored. This offers no security when all persistent data is leaked. before they are stored, to create what we call protected embeddings, such that they can still be used for $\ell_{2}$-based authentication while resisting model inversion attacks? The advantage of post-processing defenses is that they do not require re-training of the ML model, and they are agnostic to any specific model inversion attack. They offer a complementary approach to many in-training defense methods [2022:bido, 2021:mid, 2024:midre, 2023:defense-gan], which offer no formal security guarantees, nor do they aim to develop an attack-agnostic defense. Remarkably, despite their appeal, post-processing defenses have not yet been considered for tackling model inversion.
To understand why, it is perhaps useful to draw parallels with how hashing protects against the same threat model of password database breaches [2019:ml-crack-password]. It is standard practice to store passwords salted-and-hashed using standard cryptographic hash functions like SHA256, so that inversion of the plaintext passwords from a leaked password database becomes hard [1979:salt-and-hash]. However, such hash functions do not work for protecting embeddings generated from ML models. This is because ML models are designed to tolerate noise: Embeddings generated from two different face images are considered to be of the same person if their generated embedding vectors are close in $\ell_{2}$ distance. In face authentication, the user being authenticated must have an embedding vector that is close to the one expected, unlike passwords that must match exactly. However, standard cryptographic hash functions (eg. SHA256) used for salting-and-hashing passwords cannot tolerate any errors in inputs, i.e., even a single bit change in input induces a large change in output of the hash function. Hence, standard hash functions like SHA256 cannot be used to protect ML embedding vectors.
@@
Post-processing defenses for tackling model inversion attacks on modern ML models have not been demonstrated so far. Nonetheless, there have been partially-secure post-processing candidates that one could leverage from prior literature as a starting point [2011:survey-biocrypto, 2018:survey-biocrypto]. These schemes allow noisy inputs to authenticate successfully as desired. Indeed, in our evaluation, we show that these schemes are effective against prior model inversion attacks which target unprotected embeddings [2024:bob], reducing their success rate from 99% to $2\%$. However, as we will show, prior protection schemes are still breakable and do not enjoy cryptographically strong properties that are ideally desired.
@@
We evaluate our Pipe attack against two of the most potent partial protection mechanisms we know of: Facial-Fuzzy Extractor (Facial-FE) [2021:facial-fe] and Multispace Random Projection (MRP) [2007:mrp]. Pipe highlights the weaknesses of these constructions when applied to face authentication under the realistic threat model of full data leakage. Experimentally, Pipe has attack success rate of over 47% for embeddings generated from state-of-the-art models like Facenet [2015:facenet] and ArcFace [2022:arcface] against both of these partial defenses (see Tables III and IV). In most settings, the observed attack success rate is well over 89%.
Figure 1 shows a successfully recovered face from Facial-FE-protected embeddings using Pipe for a quick illustration. It is close to the original image both visually and as per image similarity metrics like LPIPS [2018:LPIPS], showing the privacy-invasive nature of the attack. The normalized embeddings of these images are also close in $\ell_{2}$ distance.
@@
Does there exist a cryptographic primitive for $\ell_{2}$-based authentication that would provably defeat all model inversion attacks, regardless of the underlying ML model, with the full-leakage threat model? Three properties would be required by such a primitive: noise tolerance, security under full leakage, and utility. One can think of it as an analogue to the salt-and-hash mechanism used for passwords, except that this primitive must tolerate noise in the input vector. Additionally, it should be hard to recover an input close to the original input even when given all outputs. We introduce this ideal primitive in Section 3.
As a key conceptual contribution (see Theorem 2), we show that the ideal primitive can be achieved if one constructs a secure fuzzy extractor—a cryptographic construct [2004:fe] that predates model inversion attacks by a decade—for the desired $\ell_{2}$ norm vector comparison. Finally, as a technical contribution, we propose a candidate construction called the L2FE-Hash, a lattice-based fuzzy extractor (FE) that enables $\ell_{2}$ norm vector comparison on hashed outputs. We formally show that it is computationally secure under certain assumptions on the input distribution (see Theorem 1). We empirically show that it has reasonable accuracy for face authentication with state-of-the-art ML models. We also confirm that L2FE-Hash resists prominent prior model inversion attacks as well as our adaptive Pipe attack in our experimental evaluation (see Section 8).
@@
While we have chosen the face authentication system as a concrete application, aligning with the original work on model inversion [2015:model-inversion], the threat of inversion attacks on embedding vectors extends well beyond that. All our security properties, attacks, and definitions are generic—they are defined abstractly for models that produce similar high-dimensional embedding vectors for similar inputs. We define closeness with respect to $\ell_{2}$ distance since it is common in face authentication, but one can easily extend our definitions to any distance metric. One can readily consider modalities that rely on $\ell_{2}$ distance, like audio and text [2016:recon-audio, 2022:text-recon-attack] in future work. Likewise, one can consider other security applications such as image similarity checking for copyright infringement or CSAM detection [2021:neural-hash].
@@
To the best of our knowledge, this is the first work that connects two long-standing but disparate lines of research, namely on model inversion attacks and on cryptographic fuzzy extractors. First, we show a new model inversion attack called Pipe which reliably defeats existing post-processing mechanisms, highlighting the lack of practical defenses with desirable cryptographic properties. Second, we describe a candidate fulfilling all ideal properties practically, called L2FE-Hash. We prove that fuzzy extractors are robust against inversion attacks generically, and show that L2FE-Hash is an $\ell_{2}$-norm fuzzy extractor construction that is empirically robust against Pipe, while having a usable accuracy for face authentication.
@@ 2 The Problem
Protecting long-term secrets, like authentication credentials, is an important goal in the event of catastrophic data breaches. We explain how this problem manifests in the context of vector embeddings used in modern ML models for face authentication, following the inspiration set in [2015:model-inversion].
@@ 2.1 Application: ML Face Authentication
Billions of users have had password credentials leaked via data breaches, posing serious threats to security and privacy. As authentication services, for example ML-based face authentication like ClearView [2023:clearview-ai] and crypto wallets [2024:biometric-wallet] store credentials (embeddings) derived from face images persistently, they face a similar risk from breach or theft [2024:outabox-breach].
@@ 2.2 Inversion of Unprotected ML Embeddings
Consider an attacker who gets unauthorized access to the database storing sensitive embedding vectors persistently. If no protection mechanism was implemented (i.e. $\mathcal{M}$ outputs were stored directly in Figure 2), then the “unprotected” embedding $x$ retains significant information about the original image, and there is a risk that the original image can be inverted from $x$ approximately. Recent prior works have shown that ML models can be trained [2023:stylegan-recontruction-attack, 2023:id3pm-diffusion-inversion-attack, 2024:bob] to perform such inversion for standard face recognition ML systems ($\mathcal{M}$) such as Facenet [2015:facenet] and ArcFace [2022:arcface]. If successful, the inverted image is close enough to the original face image that it passes authentication with $\mathcal{M}$. Prior works have reported attack success rates of over $90\%$ (see Section 6 too). Thus, highly effective inversion attacks on unprotected embeddings are known, posing a serious privacy risk.
@@ 2.3 Protecting Embeddings via Post-processing
A good defense mechanism should protect embedding vectors derived from sensitive inputs when they are persistently stored. Ideally, the defense should not be attack-specific and should generically defeat all feasible attacks. Moreover, we are not seeking solutions for securing persistent database storage via additional trust in hardware [2016:sgx-explained] or in third parties to securely store cryptographic keys.
@@
Many prior defenses are not in the post-processing regime. They are methods that work during model training to reduce the amount of sensitive information held by the resulting embeddings [2022:bido, 2023:defense-gan, 2021:mid, 2024:midre]. But such mechanisms have had limited success. Recent works show that the defended model remains susceptible to prior inversion attacks, with $\geq 30\%$ users being re-identified for models that have accuracy $\geq 80\%$ [2022:bido, 2024:midre, 2023:re-thinking]. Therefore, we leave this alternative line of mechanisms out of scope. Besides, they are complementary to post-processing mechanisms.
@@ 2.4 Threat Model
To keep attack assumptions minimal, we work in the black-box threat model. We adapt the black-box model presented in [2015:model-inversion] and assume the attacker only has query (or API) access to the target model $\mathcal{M}_{prot}$ (see Figure 2). The query exposes protected embeddings, instead of classification results as presumed in [2015:model-inversion] for generality, since ML models today are used in non-classification tasks too. Moreover, protection schemes used by the target models are publicly available algorithms known to the attacker.
We assume that the attacker has access to a public dataset of face images, similar to that assumed in the original attack [2015:model-inversion]. Our public dataset does not overlap with the leaked database, and the attacker can query $\mathcal{M}_{prot}$ on this dataset to obtain corresponding embeddings. These assumptions are practical and common to prior attacks [2024:bob, 2021:ked-mi, 2020:gmi]. This is also analogous to how password crackers use real-world password distributions gleaned from leaked password databases [2019:ml-crack-password, 2019:password-crack-survey].
@@ 3 An Ideal Post-Processing Primitive
The need for post-processing protection mechanisms has been recognized in prior cryptographic literature [2011:survey-biocrypto, 2018:survey-biocrypto], but their desired security properties have not been carefully established in the context of model inversion. In particular, no secure mechanisms from prior works satisfy the desired properties, as we will experimentally show in Section 6.
@@ II Input image obtained during enrollment
Item 3 is necessary to prevent trivial functions that compress inputs into a very small output domain. If output entropy is too small, the adversary would be able to guess inputs by brute-forcing. Moreover, it would render the primitive useless in applications that want to use the output to distinguish various inputs. For example, a primitive designed for face authentication should map input embeddings corresponding to different users to different protected embeddings. Hence, the output entropy required for Item 3 depends on the application use case and is often related to the input entropy. We use the standard computational HILL entropy [1999:extractor-universal-hash-hill] as a measure of the output entropy.
@@ 3.2 Relation to Fuzzy Extractors
We will later show formally that the desired properties are achieved by a fuzzy extractor. Briefly, the Gen and Rep subroutines of a fuzzy extractor work similarly to the Gen and Rep subroutines of our ideal primitive, albeit with different security formulations. However, existing fuzzy extractor constructions are designed for $\ell_{0}$ distances, such as the Hamming distance or set difference, to suit the use case of iris or fingerprint authentication [2004:fe]. There is no secure construction of an $\ell_{2}$-based fuzzy extractor, as needed for practical use with face distributions yet. We address this gap by proposing an ideal primitive for an $\ell_{2}$-based authentication system, and give a working candidate construction L2FE-Hash of the ideal primitive in Section 7.
@@ 4 Existing Protection Schemes
As a starting point for post-processing defenses, one can consider two cryptographic schemes that are state-of-the-art in their respective family of concepts: a Facial-Fuzzy Extractor and Multispace Random Projection. We detail their constructions and explain why they are only partially secure under the threat model of full leakage. They serve as two main targets for our new attack Pipe presented in Section 5, and experimentally confirm it in Section 6.